A man borrowed P10,000 from a bank and agreed to pay the loan at the end of 10 months. The bank
discounted the loan and gave him P9,500 in cash.
a) What was the rate of discount?
b) What was the rate of interest?
c) What was the rate of interest for one year?
1
Expert's answer
2021-09-21T02:06:38-0400


(a)

Here, P = 1000r = 0.08, and m = 1

  Thus, i = r = 0.08 and n = 3, so

 "A=P(1+r\/m)^n"

"=1000(1+0.08\/1)^3"

"=1000(1.08)^3"

=1259.71

or $1259.71.

(B)

Here, P = 1000r = 0.08, and m = 2

  Thus,  and n = (3)(2) = 6, so


 "A=P(1+r\/m)^n"

"=1000(1+0.08\/2)^3"

"=1000(1.04)^3"

  $1265.32.

(C)

Here, P = 1000r = 0.08, and m = 4.

  Thus,  and n = (3)(4) = 12, so

 "A=P(1+r\/m)^n"

=1000(1+0.08/4)^3

"=1000(1.02)^3"

  $1268.24.